# Special divisors in special divisor classes on graphs

### v2

@@ 4.3. The relation between $\mathcal{E}$- and $V$-reduced divisors
Now assume conversely that $\underline{d}$ is $V$-reduced. In particular, $\underline{d}$ is effective away from $V$. Let $\mathcal{E}$ be the $\mathbb{R}$-divisor constructed in Lemma 4.8. That is, $\mathcal{E}$ is supported on $V$, has degree $d$ and $q_{\mathcal{E}}(\underline{d})$ has value 00 on all vertices in $V$. We claim that $\underline{d}$ is $\mathcal{E}$-reduced.
@@
where the minimal value of $f$ is 00.
@@
where $-f+k\cdot\mathbb{1}_{V(G)}$ is the function obtained from $-f$ by adding $k$ to its value at each vertex. Thus $-f+k\cdot\mathbb{1}_{V(G)}$ has minimal value 00 and maximal value $k$, the latter achieved exactly on vertices in $U_{0}(f)$.
@@
$-f+k\cdot\mathbb{1}_{V(G)}=\sum_{i=l}^{k}\mathbb{1}_{F_{i}}.$
@@
The $\mathbb{R}$-divisor $\mathcal{E}$ constructed in the proof of Theorem 4.9 is $(2.5,0,3.5)$. Then $q_{\mathcal{E}}(\underline{d})$ is 00 on both $v_{1}$ and $v_{3}$ and has outgoing slope at $v_{1}$ and $v_{3}$ equal to $0.75$ along each edge.
@@ 5.1. Uniform divisors and special divisor classes
The degree of the canonical divisor is $22g-$ 2, where $g$ denotes the genus of the graph. Indeed, we have
@@
A divisor $\underline{d}$ with negative degree $d$ cannot be effective. Since the degree of the residual $\underline{d}^{*}$ is $2g-2-d$, it follows that the degree of a uniform divisor satisfies $20\leq d\leq 2g-$ 2. The same holds for the degree of a special divisor class.
@@
A special divisor class not containing a uniform representative. Consider the graph $G$ with three vertices of weight 00 and two edges between each pair of vertices (see Figure 1). The valence of each vertex is $4$ and thus the canonical divisor on $G$ is $(2,2,2)$.
@@ 5.2. Uniform representatives in special divisor classes
The upper bound for $\underline{d}_{v}$ in Theorem Theorem A, $\max\left\{\operatorname{val}_{\mathrm{wl}}(v)-1,2g_{v}-2+\operatorname{val}(v)\right\}$, is clearly weaker than the upper bound required for uniform divisors. However, the only case in which $\operatorname{val}_{\mathrm{wl}}(v)-1>2g_{v}-2+\operatorname{val}(v)$ is if $g_{v}=0$ and $\operatorname{val}_{\mathrm{wl}}(v)=\operatorname{val}(v)$; that is, if $v$ has weight zero and no adjacent loop. In this case, the upper bound in Theorem Theorem A is higher by $1$ compared to the upper bound for uniform divisors. In all other cases, the bound in the theorem coincides with that for uniform divisors. Thus if every vertex of $G$ with $g_{v}=0$ is adjacent to a loop, then every special divisor class contains a uniform representative.

### v3

@@ 4.3. The relation between $\mathcal{E}$- and $V$-reduced divisors
Now assume conversely that $\underline{d}$ is $V$-reduced. In particular, $\underline{d}$ is effective away from $V$. Let $\mathcal{E}$ be the $\mathbb{R}$-divisor constructed in Lemma 4.8. That is, $\mathcal{E}$ is supported on $V$, has degree $d$ and $q_{\mathcal{E}}(\underline{d})$ has value 0 on all vertices in $V$. We claim that $\underline{d}$ is $\mathcal{E}$-reduced.
@@
where the minimal value of $f$ is 0.
@@
where $f+k\cdot\mathbb{1}_{V(G)}$ is the function obtained from f-f by adding $k$ to its value at each vertex. Thus $f+k\cdot\mathbb{1}_{V(G)}$ has minimal value 0 and maximal value $k$, the latter achieved exactly on vertices in $U_{0}(f)$.
@@
$f+k\cdot\mathbb{1}_{V(G)}=\sum_{i=l}^{k}\mathbb{1}_{F_{i}}.$
@@
The $\mathbb{R}$-divisor $\mathcal{E}$ constructed in the proof of Theorem 4.9 is $(2.5,0,3.5)$. Then $q_{\mathcal{E}}(\underline{d})$ is 0 on both $v_{1}$ and $v_{3}$ and has outgoing slope at $v_{1}$ and $v_{3}$ equal to $0.75$ along each edge.
@@ 5.1. Uniform divisors and special divisor classes
The degree of the canonical divisor is 2g22g-2, where $g$ denotes the genus of the graph. Indeed, we have
@@
A divisor $\underline{d}$ with negative degree $d$ cannot be effective. Since the degree of the residual $\underline{d}^{*}$ is 2g2d2g-2-d, it follows that the degree of a uniform divisor satisfies $0\leq d\leq 2g-2$. The same holds for the degree of a special divisor class.
@@
A special divisor class not containing a uniform representative. Consider the graph $G$ with three vertices of weight 0 and two edges between each pair of vertices (see Figure 1). The valence of each vertex is $4$ and thus the canonical divisor on $G$ is $(2,2,2)$.
@@
Remark 5.4.
Our interest in uniform representatives comes from algebraic geometry. More precisely, results of [Chr23] establish a connection between uniform representatives and line bundles on nodal curves that satisfy the classical Clifford inequality. See [BCM25] for a detailed discussion, and, for example, [Cap94, Bak08, CLM15, Len17, CC19, Chr24] for more general context on line bundles on nodal curves.
@@ 5.2. Uniform representatives in special divisor classes
The upper bound for $\underline{d}_{v}$ in Theorem 5.6, $\max\left\{\operatorname{val}_{\mathrm{wl}}(v)-1,2g_{v}-2+\operatorname{val}(v)\right\}$, is clearly weaker than the upper bound required for uniform divisors. However, the only case in which $\operatorname{val}_{\mathrm{wl}}(v)-1>2g_{v}-2+\operatorname{val}(v)$ is if $g_{v}=0$ and $\operatorname{val}_{\mathrm{wl}}(v)=\operatorname{val}(v)$; that is, if $v$ has weight zero and no adjacent loop. In this case, the upper bound in Theorem 5.6 is higher by $1$ compared to the upper bound for uniform divisors. In all other cases, the bound in the theorem coincides with that for uniform divisors. Thus if every vertex of $G$ with $g_{v}=0$ is adjacent to a loop, then every special divisor class contains a uniform representative.
Remark 5.8.
An alternative proof of Theorem 5.6 can be given using Spencer Backman’s theory of partial graph orientations [Bac17]. We give a sketch of this construction: One may assume that $g_{v}=0$ for all vertices $V(G)$ by working on a weightless graph $G^{\prime}$ obtained from $G$ by adding at each vertex $g_{v}$ loops. Then linear equivalence on $G$ and $G^{\prime}$ coincide. Suppose first that we have $d\leq g-1$. Since $\underline{d}$ is effective, we have by [Bac17, Theorem 4.10] that $\underline{d}$ is given, up to linear equivalence, by a sourceless partial orientation. This means that $0\leq\underline{d}_{v}\leq\operatorname{val}(v)-1+g_{v}$ with equality in the upper bound only possible when there are no loops based at $v$ and $g_{v}=0$. Similarly, for $d\geq g-1$, one can use [Bac17, Algorithm 4.3] to show that the residual $\underline{d}^{*}$ can be given, up to linear equivalence, by a sinkless partial orientation. Furthermore, this partial orientation can be arranged to have no sources at vertices that are adjacent to a loop edge. It follows that $1\leq\underline{d}^{*}_{v}\leq\operatorname{val}(v)-2+g_{v}$ where the lower bound can only be achieved at weightless vertices not adjacent to any loop edges. This in turn gives $g_{v}\leq\underline{d}_{v}\leq 2g_{v}-1+\operatorname{val}(v)$, where equality in the upper bound again is only possible when there are no loops based at $v$ and $g_{v}=0$.